# Shifting Perspectives: Steering Vector Ensembles for Robust Bias Mitigation in LLMs

### v1

@@ 5 Results
In this section, we present a comprehensive evaluation of our bias mitigation methods across three instruction-tuned models: Mistral, Llama, and Qwen. We first assess the impact of individually optimized steering vectors (ISVs) on bias reduction using the Bias Benchmark for QA (BBQ) and on general language performance using MMLU. Next, we compare these results to Steering Vector Ensembles (SVEs), which average multiple ISVs to capture a more universal bias representation. Finally, we analyze the interplay between bias mitigation and general performance, and evaluate the out-of-distribution generalizability on unseen intersectional bias axes.
5.1 Effectiveness of Individual Steering Vectors
Our results show that individually tuned steering vectors, denoted as ISV in Table 2, significantly improve bias mitigation across all three models. As shown in Table 3, ISVs yield average improvements of 12.2% in BBQ accuracy for Mistral, 4.73% for Llama, and 3.20% for Qwen relative to their respective baselines. These results align with prior work in AI safety, such as toxicity reduction in Wang et al. (2024a) and Turner et al. (2024).
Building on these insights, we evaluate Steering Vector Ensembles (SVE), which combine multiple individual steering vectors via averaging. In Table 2, we observe that in many cases, though not all, SVE outperforms individually tuned steering vectors on the axis they have been optimized on, highlighting its effectiveness as a method.
Model Steering Vector BBQ MMLU
Mistral Baseline 53.6 50.3
Average ISV 65.5 42.8
Merged Datasets 40.5 30.5
SVE 69.3 46.6
Llama Baseline 75.9 52.9
Average ISV 80.1 56.0
Merged Datasets 69.5 42.7
SVE 81.6 58.1
Qwen Baseline 84.7 66.7
Average ISV 86.1 66.8
Merged Datasets 85.8 66.7
SVE 86.9 66.9
Table 3: Comparison of performance on BBQ and MMLU across three models. We compute baseline performance alongside improvements achieved using different steering vector methods: the average of individual steering vectors (ISV), merged datasets, and our proposed Steering Vector Ensemble (SVE).
Figure 4: Accuracy versus Steering Vector Coefficient for the Mistral, Llama and Qwen models on BBQ and MMLU. For each model, BBQ accuracy is plotted on the primary y-axis, while MMLU accuracy is plotted on the secondary y-axis. Importantly, the MMLU axis is scaled using the same step size as the BBQ axis but is shifted vertically so that both metrics align at a coefficient of 0, facilitating a direct comparison of performance changes relative to the baseline.
5.2 Steering Vector Ensembles (SVE) Outperform Other Methods
In Table 3, we compare various baselines on the full BBQ dataset and MMLU. We compute accuracy for BBQ and MMLU for each of the nine individual steering vectors, and take the average score (Average ISV). While BBQ scores improved, MMLU performance varied across models: applying individual steering vectors led to a 7.5% decrease in MMLU accuracy for Mistral but a 5.2% increase in Llama, and remained similar for Mistral, highlighting a potential trade-off between fairness and general capabilities that varies by architecture.
Additionally, we investigate whether simply aggregating all contrastive pairs across nine bias axes into a single dataset has a similar effect to averaging the steering vectors themselves. We create a steering vector from this single large contrastive dataset, named Merged Datasets in Table 3. We observe performance below individual steering vectors for both BBQ and MMLU in all three models, and significantly below the baseline performance in Mistral and Llama. This result suggests that highly specialized, targeted contrastive datasets are more effective than a one-size-fits-all approach, likely because overly general datasets fail to capture distinct patterns, leading to weaker learned representations. Thus, an alternative method of combining vectors without dataset merging, such as SVE, is necessary.
We observe in Table 3 that SVEs outperform all other methods on both BBQ and MMLU in all cases, with the sole exception of MMLU on Mistral. These results support our hypothesis outlined in Section 3.3, validating the idea that averaging across multiple bias concepts reduces variations unrelated to bias, which reinforces a more generalized bias representation and mitigates the dataset dependency issues that prevent generalization, as discussed in Tan et al. (2024).
5.3 Relationship between BBQ and MMLU
We examine how bias mitigation, quantified via BBQ accuracy, and general language performance, measured by MMLU accuracy, vary as a function of the steering vector coefficient. In our experiments, the coefficient spans from -5 to 5, with 0 representing the baseline result (i.e., no steering vector intervention). To facilitate a direct comparison between the two metrics, we scale the MMLU axis using the same step size as the BBQ axis and shift it vertically so that both metrics align at a coefficient of 0.
Figure 4 shows that for the Mistral model, increasing the coefficient from 0 to 5 results in an improvement in BBQ accuracy from 53.6% to 69.5%, while MMLU accuracy declines from 50.1% to 46.0%. In contrast, the Llama and Qwen exhibit more balanced responses, where MMLU remains stable as BBQ accuracy increases.
These trends indicate that while steering vectors can effectively enhance bias mitigation (as reflected by improved BBQ scores), their influence on general model performance is model-dependent. For instance, stronger models such as Qwen, which already demonstrate high baseline performance, exhibit minimal variability in MMLU scores across different coefficients, suggesting that steering vector interventions may become more effective as models scale. Overall, these findings underscore the importance of carefully calibrating the steering vector coefficient for each model.
BBQ Axis Mistral Llama Qwen
R × G R × SES R × G R × SES R × G R × SES
Baseline 55.0 55.7 80.0 83.3 86.6 89.2
Age 64.6 68.3 81.8 84.6 89.7 90.1
Appearance 66.3 66.4 80.9 83.3 86.5 87.8
Disability 71.7 68.2 86.5 86.2 89.8 90.7
Gender 70.8 68.4 87.4 83.0 87.9 88.3
Nationality 70.5 69.2 86.7 83.6 90.2 90.5
Race 68.4 62.3 84.4 84.3 90.2 90.6
Religion 64.6 68.3 87.5 86.5 86.3 87.1
Sexuality 62.6 66.0 86.3 87.7 85.7 87.9
Socioeconomic 63.5 65.6 87.5 86.3 87.9 90.0
SVE 64.5 68.3 87.3 86.9 89.3 90.2
Table 4: Baseline, 9 ISV, and SVE accuracies for Race × Gender and Race × Socioeconomic bias axes in Mistral, Llama, and Qwen, shown as percentages. Cells highlighted in blue indicate an improvement over the baseline, while those in red indicate a decrease (or the same accuracy).
5.4 Generalization and Robustness
To assess the robustness of our steering vector methods, we evaluate whether vectors optimized on one bias axis generalize to intersectional bias domains that were not used during training. Table 4 presents the accuracies for two intersectional tasks, Race × Gender and Race × Socioeconomic, across Mistral, Llama, and Qwen. These intersectional axes serve as out-of-distribution test cases.
Our results show 5 out of 9 individual steering vectors, as well as the SVE outperform the baseline, further supporting our hypothesis that SVE will demonstrate a more stable performance across both in-distribution and out-of-distribution settings.
@@ 6 Conclusion
In this work, we applied steering vectors to bias mitigation in large language models and evaluated multiple approaches across three models. Our experiments show that individually optimized steering vectors led to significant improvements in BBQ accuracy. Our use of Bayesian optimization enabled us to systematically identify effective contrastive datasets across nine bias axes, further refining the tuning of individual steering vectors.
Building on these findings, we explored the cumulative effects of combining multiple steering vectors and introduced Steering Vector Ensembles (SVE) as a generalizable and efficient strategy for fairness interventions. We further analyzed the impact of these interventions on overall model performance using the MMLU benchmark, revealing that the effect on performance varies across models. Overall, our results demonstrate that SVE not only enhances bias mitigation compared to individual steering vectors but also provides a more robust and generalized intervention, with promising implications for improving fairness and safety in large language models.
@@ 6.1 Future Work
Contrastive Datasets
Although our work relied on a uniform dataset format with variations in text content, alternative contrastive dataset structures such as those shown in Zou et al. (2023), and Rimsky et al. (2024) could be applied. In addition, extending Bayesian optimization to include the selection of layers for intervention, optimizing based on accuracy improvements, represents a promising direction.
Steering Vectors
While we focus on BBQ and MMLU, future studies could expand the evaluation of steering vectors by employing additional benchmarks. This broader evaluation could help address current limitations and validate the generalizability of our approach.
SVEs
While Steering Vector Ensembles (SVE) have shown promising improvements over individual steering vectors, further work is needed to determine the optimal combination of individual steering vectors. Future research should explore whether different subsets of steering vectors yield more effective ensembles and consider alternative aggregation methods such as weighted averages or the median vector, which may be less susceptible to outliers. Moreover, applying SVEs to additional domains beyond bias mitigation in language models will help the broader utility of this approach.
@@ 7 Limitations
Our experiments were conducted on 7B and 8B parameter models, which may not fully capture emergent abilities related to bias observed in larger models, such as moral self-correction that tends to emerge in models with 22B parameters or more, as noted in Ganguli et al. (2023). Due to computational constraints, we were unable to evaluate such larger models.
Our MMLU results suggest that steering vectors have less impact on higher-performing models, however, MMLU may not capture all aspects of language understanding and reasoning. Incorporating additional benchmarks, such as GLUE Wang et al. (2018) and HellaSwag Zellers et al. (2019), would provide a more complete assessment of the broader effects of steering vector interventions.
@@ Acknowledgements
Appendix A Additional Analysis
Figure 5: The full cosine similarity matrix over all the hidden layers for the 9 BBQ steering vectors for Mistral.
Figure 6: The full cosine similarity matrix over all the hidden layers for the 9 BBQ steering vectors for Llama.
Figure 7: The full cosine similarity matrix over all the hidden layers for the 9 BBQ steering vectors for Qwen.
Figure 8: A clustering in the similarities of the steering vectors for the 9 BBQ axes can be observed for later layers and layers that are closer together for Mistral. The layer at which the largest cluster appears is dataset dependent e.g. hidden layer 19 for the age axis and layer 15 for the socioeconomic axis.

### v2

@@ 5 Results
In this section, we present a comprehensive evaluation of our four bias mitigation methods across four evaluation datasets. Table 2 reports the performance of the four bias mitigation strategies: Prompting, Self-Debias, parameter-efficient fine-tuning (Finetuned), and steering vectors (Steering Vec.), as well as no intervention (Baseline) using the Mistral 7B model. We consider nine social axes on the Bias Benchmark for QA (BBQ), three axes on StereoSet (ICAT score), six axes on CLEAR-Bias, and general model performance via MMLU. Below we walk through the major findings for each dataset, followed by general patterns and an analysis of the interplay between the dataset specific findings.
BBQ.
On the BBQ test set of 600 examples per axis, steering vectors consistently outperformed all other methods on eight of nine axes. We expect to see a larger gain over the baseline here in both the finetuned model and steering vectors as they were pretrained on a BBQ training dataset, however, finetuning has the third highest gain on average (6.2%), behind both prompting (8.9%) and steering vectors (12.8%), though finetuning still retains the highest accuracy for religion.
StereoSet.
We next evaluate on both the intrasentence and intersentence tasks of StereoSet, reporting the ICAT score (higher = less stereotype). Here fine-tuning exhibits the strongest performance across all three axes, followed by steering vectors and then the baseline, with the exception of religion, where the religion steering vector under-performs on both the BBQ and StereoSet dataset. Prompting under-performs on all axes in this task; we posit that the low performance of prompting stems from the nature of perplexity based evaluations, i.e. a prompt mentioning bias is more likely to occur before a biased sentence than an unbiased one. Note that Self-Debias is not applicable on this task as the method is not designed to work with perplexity based evaluations.
CLEAR-Bias.
CLEAR-Bias measures adversarial robustness using known jailbreak prompts. As mentioned in Section 1, fine-tuning remains vulnerable to adversarial attacks Zhan et al. (2024), and risk false alignment, where models merely mimic patterns of their finetuning data without truly understanding task requirements Kung and Peng (2023); Wang et al. (2024). Finetuning is outperformed by prompting on 3 out of 6 axes, and by steering vectors on 5 out of 6 axes, which suggests that the LoRA adapters did not converge on a robust, bias-averse subspace, whereas steering vectors applied at inference without weight updates, more reliably mitigate stereotype activation under adversarial conditions.
MMLU.
Finally, we use MMLU as a proxy to measure general model performance and assess the collateral impact of each bias mitigation method. Here the baseline model achieves 50.7 % accuracy, and bias mitigation methods negatively impact this. Steering vectors reduce this by only a small amount (3.9%), suggesting that it is the least disruptive bias mitigation method tested as it incurs the smallest trade-off between bias mitigation and overall task performance. In contrast, finetuning decreases MMLU performance by 23.4%, suggesting the finetuned model has largely overfitted to the BBQ training dataset.
In summary, steering vectors deliver the strongest and most consistent bias reductions on targeted QA tasks (BBQ, CLEAR-Bias), with only modest impact on general capabilities (MMLU). Parameter-efficient fine-tuning still excels on StereoSet, but at the cost of larger performance degradation elsewhere as a result of overfitting. Prompting and Self-Debias provide lightweight interventions but yield smaller and less reliable improvements on bias tasks whilst still incurring a larger MMLU trade off than steering vectors. These results demonstrate that activation steering offers a compelling, computationally efficient, and broadly applicable mechanism for bias mitigation in large language models.
@@ 6 Conclusion
In this work, we apply steering vectors to bias mitigation and determine whether the method can be applied to unseen datasets. Our experiments show that steering vectors consistently outperform three other bias mitigation methods across the BBQ and CLEAR-Bias datasets, achieving an average accuracy gain of 12.6%. Steering vectors also have the lowest impact on MMLU performance (-3.9%), in comparison to finetuning which showed the largest degradation in model performance (-23.4%). While steering vectors still showing an improvement over the baseline on the perplexity based StereoSet evalutaion, they underperform compared to finetuning.
By measuring linear separability using a two component PCA and a Logistic Regression classifier, we are able to identify the optimal layer to intervene on for each steering vector, confirmed with a further layer-by-layer validation accuracy task. We continue by tuning the steering coefficient, in order to find a steering vector setup that will generalize across datasets.
By applying steering vectors at inference time, without modifying any model weights, we deliver a plug-and-play intervention that is both interpretable and computationally lightweight, and achieves substantial bias reduction with minimal impact on core performance, offering a practical path toward fairer and safer LLM deployments.
@@ 6.1 Future Work
Contrastive Datasets.
We use a contrastive dataset structure based on those shown in Zou et al. (2023), and Rimsky et al. (2024), however, other setups such as varying tokens such as he/she for gender based bias mitigation or words that reinforce or contrast a concept may lead to alternative findings.
Multi-dimensional steering.
Rather than a single principal component, future work could explore controlling along multiple PCA axes simultaneously, enabling finer-grained adjustments and potentially uncovering subtler bias facets.
Cross-axis interactions.
Investigate whether combining or orthogonalizing bias vectors across different social dimensions (e.g. gender vs. race) yields synergistic effects or mitigates unintended cross-bias amplification.
Adaptive coefficient selection.
Develop automated strategies, such as validation-based or reinforcement-learning controllers, to dynamically adjust steering strength per input, which could allow optimization of the bias vs general model performance trade-off in real time.
Broader safety applications.
Apply steering vectors to other forms of harmful behaviors (e.g. toxicity, misinformation) and assessing real-world impact in downstream tasks, for example, in social media data.
Overall, our findings underscore the potential of representation-level interventions as a lightweight yet effective complement to existing debiasing paradigms, pointing the way toward more robust and generalizable fairness safeguards in future LLM deployments.
@@ 7 Limitations
Our experiments were conducted on a 7B parameter model, which may not fully capture emergent abilities related to bias observed in larger models, such as moral self-correction that tends to emerge in models with 22B parameters or more, as noted in Ganguli et al. (2023). Due to computational constraints, we were unable to evaluate such larger models.
Our MMLU results suggest that steering vectors have less impact than other bias mitigation methods on general model performance, however, MMLU may not capture all aspects of language understanding and reasoning. Incorporating additional benchmarks, such as GLUE Wang et al. (2018) and HellaSwag Zellers et al. (2019), would provide a more complete assessment of the broader effects of steering vector interventions.
@@ Acknowledgements
Appendix A Layer-wise Linear Separability
Figure 5: Two component PCA graphs over all the hidden layers for the the nationality vector, with the logistic regression classifier accuracy, demonstrating the linear separability at each layer.